andrew made a trip to his cabin on the lake and back. the trip there took 3 hours and the trip back took 5 hours. he averaged 27 mph on the return trip. find the average speed on the trip there

Answer by ankor@dixie-net.com(22740)

Andrew made a trip to his cabin on the lake and back.
the trip there took 3 hours and the trip back took 5 hours.
he averaged 27 mph on the return trip.
find the average speed on the trip there
:
let s = av speed there
:
The trips there and back are equal, write a dist equation; dist = speed * time
3s = 27(5)
3s = 135
s = 135/3
s = 45 mph there
:
See if that checks out find the distances, should be equal
27*5 = 135
45*3 = 135